OpenStudy (anonymous):

If f(x) = 2x^2 − 30, find f(4).

OpenStudy (anonymous):

Basically, f(4) means that here x = 4 So: Put x = 4 in the function: \[f(4) = 2(4)^2 - 30\] Now can you find its value ??

OpenStudy (anonymous):

The second2 isanxponent.

OpenStudy (anonymous):

Yes I understood that. That is why I put 2 as power of 4..

OpenStudy (anonymous):

I said to find the value..

OpenStudy (anonymous):

I am doing flvs online school and i have to basicly teach myself, i dont understand it at all

OpenStudy (anonymous):

What is \(4^2\) = ??

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

nooooo nvm 16

OpenStudy (anonymous):

\[4^2 = 4 \times 4 = ??\] Basically power means multiply the base power times..

OpenStudy (anonymous):

Yes: \(2(16)\) = ??

OpenStudy (anonymous):

Bracket shows here nothing but multiplication; Okay in other way you can write it as: \(2 \times 16\) = ??

OpenStudy (anonymous):

32

OpenStudy (anonymous):

Now last step: \(32 - 30\) = ??

OpenStudy (anonymous):

That will be the value for f(4).

OpenStudy (anonymous):

so the answer is 2